# Towards a Unified Language Model for Knowledge-Intensive Tasks Utilizing External Corpus

### v1

@@ Abstract.
The advent of large language models (LLMs) has showcased their efficacy across various domains, yet they often hallucinate, especially in knowledge-intensive tasks that require external knowledge sources. To improve factual accuracy of language models, retrieval-augmented generation (RAG) has emerged as a popular solution. However, traditional retrieval modules often rely on large-scale document indexes, which can be disconnected from generative tasks. Through generative retrieval (GR) approach, language models can achieve superior retrieval performance by directly generating relevant document identifiers (DocIDs). However, the relationship between GR and downstream tasks, as well as the potential of LLMs in GR, remains unexplored. In this paper, we present a unified language model that utilizes external corpus to handle various knowledge-intensive tasks by seamlessly integrating generative retrieval, closed-book generation, and RAG. In order to achieve effective retrieval and generation through a unified continuous decoding process, we introduce the following mechanisms: (1) a ranking-oriented DocID decoding strategy, which improves ranking ability by directly learning from a DocID ranking list; (2) a continuous generation strategy to facilitate effective and efficient RAG; (3) well-designed auxiliary DocID understanding tasks to enhance the model’s comprehension of DocIDs and their relevance to downstream tasks. Our approach is evaluated on the widely used KILT benchmark using two variants of backbone models: an encoder-decoder T5 model and a decoder-only LLM, Llama2. Experimental results showcase the superior performance of our models in both retrieval and downstream knowledge-intensive tasks.
@@
doi: XXXXXXX.XXXXXXXccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Traditionally, the retrieval component follows an index-centric framework (Robertson and Zaragoza, 2009; Karpukhin et al., 2020; Xiong et al., 2020; Khattab and Zaharia, 2020; Zhang et al., 2023c). Although this approach is widely used, it has several drawbacks. Notably, during inference, the requirement for a large document index to search the entire corpus results in considerable memory usage and computational demands (Chen et al., 2022). Furthermore, during training, the disparate model structures for retrieval (matching similarity) and generation (auto-regressive) (Lewis et al., 2020b; Borgeaud et al., 2022; Nakano et al., 2021; Liu et al., 2023a) hinders the joint optimization of both models and limits the understanding of the relationship between both tasks.
@@
To summarize, we make the following main contributions:
(1) Unified Language Model. We develop a unified language model capable of utilizing the knowledge of corpus for generative retrieval, closed-book generation, and retrieval-augmented generation to facilitate knowledge-intensive tasks.
(2) Ranking-oriented DocID List Generation Strategy. We surpass the current generative retrieval models that are limited to DocID generation, and propose to retrieve by generating a DocID ranking list through greedy decoding. This approach enhances the ranking quality and provides a unified solution for retrieval and generation.
(3) RAG-oriented Continuous Generation Strategy. For RAG, we introduce a continuous decoding method that first decodes relevant DocIDs, then decodes fine-grained references from the retrieved documents, and finally decodes the response.
(4) Unsupervised DocID Understanding Tasks. We design unsupervised DocID understanding tasks to deepen the model’s comprehension of the meaning behind DocIDs. This further enhances the generation quality of CorpusLM.
@@ 2.3.1. Training Data Construction
The training process involves learning from a ranked list of relevant DocIDs. However, the challenge lies in the fact that training queries are typically associated with just one or a few relevant documents, which limits the model’s capacity to learn effective ranking. To overcome this, we enhance the training data with additional relevant documents. We start with the original training queries and their respective answers. We first use a BM25 (Robertson and Zaragoza, 2009) retriever to produce an initial list of potentially relevant documents. Following this, we employ a dense re-ranker SimLM (Wang et al., 2023a) to re-rank the retrieved documents. We create a final list of top-k DocIDs by taking the reranked documents and appending them to the original labeled DocIDs, ensuring there are no duplicates in the combined list. Through this method, each query in the training set is associated with a comprehensive list of top-k relevant DocIDs, enabling the model to understand the relevance between multiple ranked DocIDs and the given query, thus reducing training confusion and improving its ranking capability. The final training data format is depicted in Figure 2, where all of the relevant DocIDs is formatted in a single natural language sequence.
@@ 2.3.2. Training Objective
where $T_{d}$ denotes the length of tokens in the sequence comprised of the target ranked DocID list. The ranking loss is designed to encourage the model to predict the correct sequence of ranked DocIDs by maximizing the log-likelihood of each token $s_{i}$, given the preceding tokens $s_{<i}$, and the query $q$. This approach helps to ensure that the model can generate a sequence of DocIDs that reflects the correct ranking order, also enable ranking-aware generative retrieval in natural language expressions, ultimately improving the retrieval performance of CorpusLM and unification with other generative NLP tasks.
@@ 2.3.3. Inference Constraints
During inference, we employ a DocID prefix tree constructed from the documents in the corpus to add constraints that ensure the generation of valid, non-repetitive DocIDs. As depicted in Figure 2(a), the process involves dynamic activation of constraints during the generation of each token: (1) If the token is the DocID start symbol ”<docid>”, we enforce prefix constraints on the succeeding sequence. (2) If the token is the DocID end symbol ”</docid>”, we lift the constraints and remove the previously generated DocID from the prefix tree to prevent duplication in subsequent generations. (3) For all other tokens, we scan backward; if ”</docid>” is encountered, scanning exits. If ”<docid>” is found, constraints are applied to the sequence following to ensure the current token’s validity within the prefix tree.
@@ 2.5. RAG: Continuous Generation Strategy
where $T_{r}$ is the length of the reference $r$, and $r_{i}$ denotes the $i$-th token in the reference sequence.
@@ 2.6. Unsupervised DocID Understanding Tasks
Table 2. Overall retrieval performance on KILT dev set. We report passage-level R-Precision (%). Models are emphasized with the best in bold and the second in underline. The symbol $\dagger$” signifies that our model achieved superior results among all baselines in a statistically significant manner (t-test, $p<0.05$).
@@ 3.2. Evaluation Metrics
For retrieval evaluation, consistent with previous works (Petroni et al., 2021; Cao et al., 2021; Bevilacqua et al., 2022; Maillard et al., 2021a; Lewis et al., 2020b) and the official KILT evaluation metrics, we use R-precision (Petroni et al., 2021) as the evaluation metric, which is by the fraction $\frac{r}{R}$, with $R$ being the number of contexts in the provenance set, and $r$ is the number of relevant contexts within the top-$R$ retrieved contexts.
For downstream evaluation, we adopt specific metrics for different downstream tasks. Specifically, as suggested in the KILT resource paper (Petroni et al., 2021), we use Accuracy for FEV, AY2, WnWi, WnCw, T-REx and zsRE; Exact Match (EM) for NQ, TQA and HoPo; ROUGE-L for ELI5; and F1 for WoW. To ensure a fair comparison with non-finetuned generators, we include the percentage of outputs containing gold answers (”has_answer”), in line with (Hoshi et al., 2023).
@@ 3.3.1. Baselines for Retrieval Tasks
Retrieval tasks are divided into two main approaches: Sparse&Dense Retrieval and Generative Retrieval. For Sparse&Dense Retrieval, we employ the following models: BM25 (Robertson and Zaragoza, 2009), a classic sparse retrieval model; DPR (Karpukhin et al., 2020) and its multi-task variant MT-DPR (Maillard et al., 2021a) for dense passage retrieval; RAG (Lewis et al., 2020b), which combines dense retrieval with seq2seq models for an enhanced generation; E5 (Wang et al., 2022b), a state-of-the-art text embedding model; and SimLM (Wang et al., 2023a), a dense passage retriever with effective pre-training methods. The Generative Retrieval methods comprise T5 (Raffel et al., 2020), a pre-trained encoder-decoder model for multi-task learning, BART (Lewis et al., 2020a), a denoising autoencoder for text generation; SEAL (Bevilacqua et al., 2022), generating sub-strings as document identifiers; CorpusBrain (Chen et al., 2022), incorporating pre-training strategies for KILT retrieval task; and Llama2 (Touvron et al., 2023b), an open-source pre-trained LLM.
The baseline retrieval models are finetuned with labeled retrieval data from KILT datasets. Except the DPR model is finetuned on each specific dataset, other dense and generative retrieval models are multi-task finetuned using retrieval data across all datasets, as multi-task training in KILT leads to improved performance (Maillard et al., 2021a; Chen et al., 2022).
@@ 3.3.2. Baselines for Downstream Tasks
For downstream tasks, we consider two primary streams: Closed-book Generation and Retrieval-augmented Generation. The Closed-book Generation models include T5 (Raffel et al., 2020), BART (Lewis et al., 2020a), and the open-source LLM named Llama2-70B (Touvron et al., 2023b), all leveraging their pre-trained parameters for generative tasks. For Retrieval-augmented Generation, we consider the following models: (1) DPR+BART (Petroni et al., 2021), combining dense passage retrieval with generative BART model; (2) RAG (Lewis et al., 2020b), an end-to-end retriever-generator model; (3) MT-DPR+FID (Maillard et al., 2021a), integrating multi-task DPR retriever with a fusion-in-decoder generator (Izacard and Grave, 2020); (4) BM25+Llama2-70B and E5+Llama2-13B&70B models reported by (Hoshi et al., 2023) that utilize Llama2 (Touvron et al., 2023b) for answer generation.
Table 3. Overall downstream performance on KILT dev set, considering both closed-book and RAG settings. In different settings, models are emphasized with the best in bold and the second in underline. ${}^{\diamondsuit}$ and ${}^{\heartsuit}$ indicates results from (Petroni et al., 2021) and (Hoshi et al., 2023), respectively. Following (Hoshi et al., 2023), the percentage of outputs containing gold answers (”has_answer”) is noted in parentheses.
@@ 3.4. Implementation Details
In our experiments, we utilize the T5-Base (Raffel et al., 2020) (T5) and Llama2-7B-Chat (Touvron et al., 2023b) (Llama2) as our backbone models with pre-trained parameters from Hugging Face (Wolf et al., 2019). During training, we set the coefficients $\lambda_{1-4}$ to 1 and the noise sampling probability $\tau$ to 0.2. Our training process involves a batch size of 512 and a learning rate of 3e-4. We use 2000 warm-up steps for the learning rate. To improve the training efficiency of the CorpusLM (Llama2) model, we employ QLoRA (Dettmers et al., 2023) and DeepSpeed (Rasley et al., 2020; Rajbhandari et al., 2020) technologies. We incorporate a maximum of top-10 and top-3 ranked passages for training the DocID decoding and answer decoding, respectively. Each unsupervised DocID understanding task consists of approximately 300k training pairs. During inference, we use dynamic constrained greedy decoding for generative retrieval tasks and greedy decoding for downstream tasks. For retrieval tasks, we limit the number of generated DocIDs to a maximum of 10. For RAG tasks, we retrieve and utilize 3 passages as the contexts for downstream tasks. All experiments are conducted on 8 NVIDIA Tesla A100 40GB GPUs.
@@ 4.1. Retrieval Performance
[…]
89.81 87.09 50.52 49.77 80.68 70.34 53.39 40.96 70.94 22.13 16.65
   w/o Decode Reference
87.21 84.17 49.09 47.97 77.95 67.92 50.47 36.42 68.50 21.16 15.85
   w/o Noise Sampling
88.31 86.48 49.41 48.53 79.14 68.90 52.78 38.81 70.29 20.17 15.05
   w/o Pipeline Decode
89.18 86.12 48.38 47.07 78.51 67.83 52.56 39.06 69.82 22.39 16.20
[…]
90.22 85.03 56.54 50.32 81.57 72.79 55.38 42.23 72.43 23.46 16.96
   w/o Decode Reference
87.40 84.49 55.14 47.89 79.61 70.53 51.88 38.25 69.58 21.61 15.70
   w/o Noise Sampling
88.04 85.43 55.42 48.65 79.77 71.51 54.25 39.66 71.79 21.31 15.66
@@ 4.2. Downstream Generation Performance
In the RAG setting, our CorpusLM models also take the lead, with CorpusLM (Llama2) demonstrating the most substantial improvements. Among small language models, our T5-based CorpusLM outperforms RAG by 12.52%, 47.98%, and 37.37% on the AY2, HoPo, and ELI5 datasets respectively. It even exceeds MT-DPR+FID, which utilizes a stronger fusion-in-decoder generator, by 9.17%, 12.21%, and 40.36% on the same datasets. Among large language models, our Llama2-based CorpusLM surpasses E5+Llama2-13B by 36.08%, 129.64%, and 37.92% on the FEVER, zsRE, and WoW datasets respectively. It even outperforms E5+Llama2-70B by 80.80%, 107.99%, and 28.51% on these datasets. Even when using the has_answer metric for E5+Llama2-70B, CorpusLM (Llama2) still achieves superior performance with gains of 1.83%, 68.51%, and 17.96% on the FEVER, zsRE, and HoPo datasets respectively. The above results validate the superior performance of our proposed multi-task learning framework and the continuous RAG strategy.

### v2

# CorpusLM: Towards a Unified Language Model on Corpus for Knowledge-Intensive Tasks

@@ Abstract.
Large language models (LLMs) have gained significant attention in various fields but prone to hallucination, especially in knowledge-intensive (KI) tasks. To address this, retrieval-augmented generation (RAG) has emerged as a popular solution to enhance factual accuracy. However, traditional retrieval modules often rely on large document index and disconnect with generative tasks. With the advent of generative retrieval (GR), language models can retrieve by directly generating document identifiers (DocIDs), offering superior performance in retrieval tasks. However, the potential relationship between GR and downstream tasks remains unexplored. In this paper, we propose CorpusLM, a unified language model that leverages external corpus to tackle various knowledge-intensive tasks by integrating generative retrieval, closed-book generation, and RAG through a unified greedy decoding process. We design the following mechanisms to facilitate effective retrieval and generation, and improve the end-to-end effectiveness of KI tasks: (1) We develop a ranking-oriented DocID list generation strategy, which refines GR by directly learning from a DocID ranking list, to improve retrieval quality. (2) We design a continuous DocIDs-References-Answer generation strategy, which facilitates effective and efficient RAG. (3) We employ well-designed unsupervised DocID understanding tasks, to comprehend DocID semantics and their relevance to downstream tasks. We evaluate our approach on the widely used KILT benchmark with two variants of backbone models, i.e., T5 and Llama2. Experimental results demonstrate the superior performance of our models in both retrieval and downstream tasks.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657778isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and ranking
CorpusLM: Towards a Unified Language Model on Corpus for Knowledge-Intensive Tasks
@@ 1. Introduction
Traditionally, the retrieval component follows an index-centric framework (Robertson and Zaragoza, 2009; Karpukhin et al., 2020; Xiong et al., 2020; Khattab and Zaharia, 2020; Zhang et al., 2023c). Although this approach is widely used, it has several drawbacks. Notably, the requirement for a large document index to search the entire corpus results in considerable memory footprint. Furthermore, during training, the disparate model structures for retrieval (matching similarity) and generation (auto-regressive) (Lewis et al., 2020b; Borgeaud et al., 2022; Nakano et al., 2021; Liu et al., 2023a) hinder the joint optimization of both models. This limitation restricts the understanding of the relationship between both tasks.
@@
2. Related Work
To enhance the performance of RAG models, various strategies have been developed, which involve training the retriever and reader modules separately using dense retrieval methods, such as DPR (Karpukhin et al., 2020), MT-DPR (Maillard et al., 2021a), DRQA (Chen et al., 2017) and FID (Izacard and Grave, 2020). Other strategies focus on joint training of both modules, either by updating the query encoder while keeping the document index static (Lewis et al., 2020b), or by updating both the query and document encoders and asymptotically updating the document index during training (Singh et al., 2021; Sachan et al., 2021; Izacard et al., 2023). Based on generative retrieval methods, SEAL (Bevilacqua et al., 2022), CorpusBrain (Chen et al., 2022) and UGR (Chen et al., 2023) train the GR models and also utilize separate generators to accomplish downstream KILT tasks. Recently, UniGen (Li et al., 2023c) introduced a unified framework for concurrent learning of GR and QA tasks. However, it uses separate decoders to generate DocIDs and answers, restricting its ability to generalize and scale.
@@ 3.3.1. Training Data Construction
The training process involves learning from a ranked list of relevant DocIDs. However, the challenge lies in the fact that training queries are typically associated with just one or a few relevant documents, which limits the model’s capacity to learn effective ranking. To overcome this, we enhance the training data with additional relevant documents. We start with the original training queries and their respective answers. We first use a BM25 (Robertson and Zaragoza, 2009) retriever to produce an initial list of potentially relevant documents. Following this, we employ a dense re-ranker SimLM (Wang et al., 2023a) to re-rank the retrieved documents. We create a final list of top-k DocIDs by taking the reranked documents and appending them to the original labeled DocIDs, ensuring there are no duplicates in the combined list.
Through this method, each query in the training set is associated with a comprehensive list of top-k relevant DocIDs, enabling the model to understand the relevance between multiple ranked DocIDs and the given query, thus reducing training confusion and improving its ranking capability. The final training data format is depicted in Figure 2, where all of the relevant DocIDs is formatted in a single natural language sequence.
@@ 3.3.2. Training Objective
where $T_{d}$ denotes the length of tokens in the sequence comprised of the target ranked DocID list. The ranking loss is designed to encourage the model to predict the correct sequence of ranked DocIDs by maximizing the log-likelihood of each token $s_{i}$, given the preceding tokens $s_{<i}$, and the query $q$.
This approach helps to ensure that the model can generate a sequence of DocIDs that reflects the correct ranking order, also enable ranking-aware generative retrieval in natural language expressions, ultimately improving the retrieval performance of CorpusLM and unification with other generative NLP tasks.
@@ 3.3.3. Inference Constraints
During inference, we employ a DocID prefix tree constructed from the documents in the corpus to add constraints that ensure the generation of valid, non-repetitive DocIDs. As depicted in Figure 2(a), the process involves dynamic activation of constraints during the generation of each token:
(1) If the token is the DocID start symbol "<docid>", we enforce prefix constraints on the succeeding sequence.
(2) If the token is the DocID end symbol "</docid>", we lift the constraints and remove the previously generated DocID from the prefix tree to prevent duplication in subsequent generations.
(3) For all other tokens, we scan backward; if "</docid>" is encountered, scanning exits. If "<docid>" is found, constraints are applied to the sequence following to ensure the current token’s validity within the prefix tree.
@@ 3.5. RAG: Continuous Generation Strategy
where $T_{r}$ is the length of the reference $r$, and $r_{i}$ denotes the $i$-th token in the reference sequence. This expects the model to learn to extract fine-grained references after the retrieved passages.
@@ 3.6. Unsupervised DocID Understanding Tasks
Table 2. Overall retrieval performance on KILT dev set. We report passage-level R-Precision (%). Models are emphasized with the best in bold and the second in underline. The symbol $\dagger$" signifies that our model achieved superior results among all baselines in a statistically significant manner (t-test, $p<0.05$).
@@ 4.2. Evaluation Metrics
For retrieval evaluation, consistent with previous works (Petroni et al., 2021; Cao et al., 2021; Bevilacqua et al., 2022; Maillard et al., 2021b; Lewis et al., 2020b) and the official KILT evaluation metrics, we use R-precision (Petroni et al., 2021) as the evaluation metric, which is by the fraction $\frac{r}{R}$, with $R$ being the number of contexts in the provenance set, and $r$ is the number of relevant contexts within the top-$R$ retrieved contexts.
For downstream evaluation, we adopt specific metrics for different downstream tasks. Specifically, as suggested in the KILT resource paper (Petroni et al., 2021), we use Accuracy for FEV, AY2, WnWi, WnCw, T-REx and zsRE; Exact Match (EM) for NQ, TQA and HoPo; ROUGE-L for ELI5; and F1 for WoW. To ensure a fair comparison with non-finetuned generators, we include the percentage of outputs containing gold answers ("has_answer"), in line with (Hoshi et al., 2023).
@@ 4.3.1. Baselines for Retrieval Tasks
Retrieval tasks are divided into two main approaches: Sparse&Dense Retrieval and Generative Retrieval. For Sparse&Dense Retrieval, we employ the following models: BM25 (Robertson and Zaragoza, 2009), a classic sparse retrieval model; DPR (Karpukhin et al., 2020) and its multi-task variant MT-DPR (Maillard et al., 2021b) for dense passage retrieval; RAG (Lewis et al., 2020b), which combines dense retrieval with seq2seq models for an enhanced generation; E5 (Wang et al., 2022b), a state-of-the-art text embedding model; and SimLM (Wang et al., 2023a), a dense passage retriever with effective pre-training methods. The Generative Retrieval methods comprise T5 (Raffel et al., 2020), a pre-trained encoder-decoder model for multi-task learning, BART (Lewis et al., 2020a), a denoising autoencoder for text generation; SEAL (Bevilacqua et al., 2022), generating sub-strings as document identifiers; CorpusBrain (Chen et al., 2022), incorporating pre-training strategies for KILT retrieval task; and Llama2 (Touvron et al., 2023b), an open-source pre-trained LLM.
The baseline retrieval models are finetuned with labeled retrieval data from KILT datasets. Except the DPR model is finetuned on each specific dataset, other dense and generative retrieval models are multi-task finetuned using retrieval data across all datasets, as multi-task training in KILT leads to improved performance (Maillard et al., 2021b; Chen et al., 2022).
@@ 4.3.2. Baselines for Downstream Tasks
For downstream tasks, we consider two primary streams: Closed-book Generation and Retrieval-augmented Generation. The Closed-book Generation models include T5 (Raffel et al., 2020), BART (Lewis et al., 2020a), and the open-source LLM named Llama2-70B (Touvron et al., 2023b), all leveraging their pre-trained parameters for generative tasks. For Retrieval-augmented Generation, we consider the following models: (1) DPR+BART (Petroni et al., 2021), combining dense passage retrieval with generative BART model; (2) RAG (Lewis et al., 2020b), an end-to-end retriever-generator model; (3) MT-DPR+FID (Maillard et al., 2021b), integrating multi-task DPR retriever with a fusion-in-decoder generator (Izacard and Grave, 2020); (4) BM25+Llama2-70B and E5+Llama2-13B&70B models reported by (Hoshi et al., 2023) that utilize Llama2 (Touvron et al., 2023b) for answer generation.
Table 3. Overall downstream performance on KILT dev set, considering both closed-book and RAG settings. In different settings, models are emphasized with the best in bold and the second in underline. and indicates results from (Petroni et al., 2021) and (Hoshi et al., 2023), respectively. Following (Hoshi et al., 2023), the percentage of outputs containing gold answers ("has_answer") is noted in parentheses.
@@ 4.4. Implementation Details
In our experiments, we utilize the T5-Base (Raffel et al., 2020) (T5) and Llama2-7B-Chat (Touvron et al., 2023b) (Llama2) as our backbone models with pre-trained parameters from Hugging Face (Wolf et al., 2019).
During training, we set the coefficients $\lambda_{1-4}$ to 1 and the noise sampling probability $\tau$ to 0.2. Our training process involves a batch size of 512 and a learning rate of 3e-4. We use 2000 warm-up steps for the learning rate. To improve the training efficiency of the CorpusLM (Llama2) model, we employ QLoRA (Dettmers et al., 2023) and DeepSpeed (Rasley et al., 2020; Rajbhandari et al., 2020) technologies. We incorporate a maximum of top-10 and top-3 ranked passages for training the DocID decoding and answer decoding, respectively. Each unsupervised DocID understanding task consists of approximately 300k training pairs.
During inference, we use dynamic constrained greedy decoding for generative retrieval tasks and greedy decoding for downstream tasks. For retrieval tasks, we limit the number of generated DocIDs to a maximum of 10. For RAG tasks, we retrieve and utilize 3 passages as the contexts for downstream tasks. All experiments are conducted on 8 NVIDIA Tesla A100 40GB GPUs.
@@ 5.1. Retrieval Performance
CorpusLM (T5) 89.81 87.09 50.52 49.77 80.68 70.34 53.39 40.96 70.94 22.13 16.65
   w/o Decode Reference 87.21 84.17 49.09 47.97 77.95 67.92 50.47 36.42 68.50 21.16 15.85
   w/o Noise Sampling 88.31 86.48 49.41 48.53 79.14 68.90 52.78 38.81 70.29 20.17 15.05
   w/o Pipeline Decode 89.18 86.12 48.38 47.07 78.51 67.83 52.56 39.06 69.82 22.39 16.20
CorpusLM (Llama2) 90.22 85.03 56.54 50.32 81.57 72.79 55.38 42.23 72.43 23.46 16.96
   w/o Decode Reference 87.40 84.49 55.14 47.89 79.61 70.53 51.88 38.25 69.58 21.61 15.70
   w/o Noise Sampling 88.04 85.43 55.42 48.65 79.77 71.51 54.25 39.66 71.79 21.31 15.66
@@ 5.2. Downstream Generation Performance
In the RAG setting, our CorpusLM models also take the lead, with CorpusLM (Llama2) demonstrating the most substantial improvements. Among small language models, our T5-based CorpusLM outperforms RAG by 12.52%, 47.98%, and 37.37% on the AY2, HoPo, and ELI5 datasets respectively. It even exceeds MT-DPR+FID, which utilizes a stronger fusion-in-decoder generator, by 9.17%, 12.21%, and 40.36% on these datasets. Among LLMs, our Llama2-based CorpusLM surpasses E5+Llama2-13B by 36.08%, 129.64%, and 37.92% on the FEVER, zsRE, and WoW datasets respectively. It even outperforms E5+Llama2-70B by 80.80%, 107.99%, and 28.51% on these datasets. The above results validate the effectiveness of our multi-task learning framework and continuous RAG strategy.